OpenStudy (anonymous):

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits.

OpenStudy (anonymous):

A square has four equal sides. Area of a square or rectangle is length times width. Just multiply two sides together and you get the area.

OpenStudy (anonymous):

Area of a square = (side)^2 In this problem, Area = (6.25)^2 = 39.0625 square feet

OpenStudy (anonymous):

length times width 6.25 times 6.25 equals 39.0625 Sq feet roundyour answer would be 39